OpenStudy (anonymous):

how do i write a question that i can answer by solving A=1000 (1+ 0.08/12)^240

ganeshie8 (ganeshie8):

compound interest ?

OpenStudy (anonymous):

what is that?

ganeshie8 (ganeshie8):

its a compound interest equation. may i ask from which topic this question is frm ?

OpenStudy (anonymous):

It could be a growth rate question.

OpenStudy (anonymous):

exponential functions

ganeshie8 (ganeshie8):

okay. compound interest is a good example of exponential growth. exponential functions grow very fast

ganeshie8 (ganeshie8):

A=1000 (1+ 0.08/12)^240 => A=1000 (1+ 0.08/12)^(12*20)

ganeshie8 (ganeshie8):

lets try framing a question

ganeshie8 (ganeshie8):

if you put 1000 $ in bank and waited 20 years, and the bank pays interest of 8 %

ganeshie8 (ganeshie8):

we have covered "1000, 0.08 and 20 " in our question. only "12" is left out, right ?

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

that 12 is simply the compounding interval. each year, the money is compounded 12 times !

ganeshie8 (ganeshie8):

the question can be, how much money the bank gives back, if you deposit $1000 for 20 years at an interest of 8% and compounded monthly

OpenStudy (anonymous):

okay got it but what about solving 2000=1000(l+r)^8

ganeshie8 (ganeshie8):

okay. id love to explain that if u pay attention :)

ganeshie8 (ganeshie8):

there it is asking to find out rate

ganeshie8 (ganeshie8):

rate of growth

OpenStudy (anonymous):

okay

OpenStudy (phi):

how to solve 2000=1000(1+r)^8? to solve divide both sides by 1000 to get 2= (1+r)^8 raise both sides to the (1/8) power to get 2^(1/8) = 1+r subtract 1 from both sides 2^(1/8)-1 = r to get a value, type 2^(1/8)-1 = in the google search window.

ganeshie8 (ganeshie8):

chikit.. you want to solve or frame a question like before ? if it is about solving then phi explanation is the best u can have